your company needs to purchase a dump truck and has narrowed the selection down to two alternatives. the 1st alternative is to purchase a new dump truck for $65,000. at the end of the seventh year the salvage value of the new dump truck is estimated to be $15,000. the 2nd alternative is to purchase a used dump truck for $50,000. at the end of 4th year the salavage of the dump truck is estimated to be $5,000. the annual profits, revenues less operation costs, are $17,000 per year for either truck. using a MARR of 18% and twenty-eight years study period, calculate the net present value for each of the dump trucka. which truck should your company purchase?


1
Expert's answer
2021-05-05T07:25:38-0400

The present value of the annual profits for either truck is determined by using USPWF:

"PAP=\$17,000\times\frac{[(1+0.18)28-1]}{0.18[1+0.18]28}=17000\times5.502=\&93,527"


The PSV for the new dump truck is determined by summing the PSVs

occurring in years 7, 14, 21, and 28. The present value for each salvage value is calculated

using SPPWF as follows:




The NPV for the purchase of the new dump truck is calculated as follows:

"NPV [New] = \$93,527 + \$6,797 + [ -\$93,822] = +\$6,502"


The PSVs for the used dump truck is determined by summing the present

value of salvage values occurring in years 4, 8, 12, 16, 20, 24, and 28. The present value for

each salvage value is calculated using SPPWF as follows:




The net present value [NPV] for the purchase of the used dump truck is calculated as

follows:

NPV = $93,527 + $5,275 + [ -$102,257] =Β -$3,455


Β The new truck has the highest NPV; therefore, your company should purchase the new

truck.